OpenStudy (anonymous):

help anyone i will fan or medal simplify 4√6+5 √6

OpenStudy (solomonzelman):

If I asked you what is 4x+5x would you tell me?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 4x+5x=? }\)

OpenStudy (anonymous):

9x

OpenStudy (solomonzelman):

yes, and when you add: \(\large\color{black}{ \displaystyle 4\sqrt{6}+5\sqrt{6} }\) you are applying the same technique

OpenStudy (solomonzelman):

when you add \(\large\color{black}{ \displaystyle 4x+5x }\) you are adding 4 x's and 5 x's (to get 9x) now, in your question this x is \(\large\color{black}{ \displaystyle \sqrt{6} }\)

OpenStudy (anonymous):

corect

OpenStudy (solomonzelman):

so, \(\large\color{black}{ \displaystyle 4\sqrt{6}+5\sqrt{6} ~=~? }\)

OpenStudy (anonymous):

so like 4+6=10 and 5+6=11 then 10+11= to my answer

OpenStudy (solomonzelman):

just like \(\large\color{black}{ \displaystyle \color{blue}{4}x+\color{blue}{5}x=\color{blue}{9} x }\) \(\large\color{black}{ \displaystyle \color{blue}{4}{\rm W}+\color{blue}{5}{\rm W}=\color{blue}{9} {\rm W} }\) \(\large\color{black}{ \displaystyle \color{blue}{4}\sqrt[7]{3}+\color{blue}{5}\sqrt[7]{3}=\color{blue}{9} \sqrt[7]{3} }\) \(\large\color{black}{ \displaystyle \color{blue}{4}\pi+\color{blue}{5}\pi=\color{blue}{9} \pi }\)

OpenStudy (solomonzelman):

when you add: \(\large\color{black}{ \displaystyle \color{blue}{4}\sqrt{6}+\color{blue}{5}\sqrt{6}~=~~~?}\)

OpenStudy (anonymous):

i think 21

OpenStudy (solomonzelman):

just add 4 and 5 , and then write sqrt(6) near it

OpenStudy (anonymous):

okay thanxs

OpenStudy (solomonzelman):

yw